A 11,000-watt radio station transmits at 880 kHz. Calculate the number of photons that are emitted per second.

________x 10 -______ photons
1
Expert's answer
2013-05-16T09:46:40-0400
P = 11000 W
ν = 880 Hz
t = 1s

Energy of one photon is
E = hν
For n photons that would be
En= nhν
Also
P = E/t = nhν/t
Thus

n = Pt/hν = (11000 [W] ×1 [s])/(6.62×10-34 [J×s]× 880× 103[s(-1)])=1.88×1031 photons